# Light4GS: Lightweight Compact 4D Gaussian Splatting Generation via Context Model

### v1

@@ 4.2 MHCM: Multiscale Hexplane compression with Context Model
Lowest-Scale Hexplane Compression: Planes in one hexplane can be divided into two groups based on their dimensions: space-only and space-time. Throughout the lowest-scale encoding process, both groups employ checkerboard modeling but use different hyperpriors based on the availability of prior information. In this approach, we designate odd positions on each plane as anchors, which are encoded with only the hyperprior. In contrast, non-anchors (even positions) are encoded using both the hyperprior and the checkerboard context that is derived by applying masked convolution on the plane anchors. Without loss of generality, we omit the superscript $l$ in this section, as it is always equal to 1. To compress the space-time group, since no prior information is available, we directly use a deep network $h_{a}()$ to extract hyperpriors as $\mathbf{z}_{c}=h_{a}(\mathbf{R}_{c})$. The hyperprior is then applied to encode space-time planes as:
@@
High-Scale Hexplane Compression: High-scale hexplanes often exhibit significant redundancy with their lower-scale counterparts. Relying solely on the checkerboard model for compression overlooks these inter-scale redundancies, as it captures only intra-scale correlations. To address this issue, we introduce an inter-scale context that encodes higher-scale planes in a coarse-to-fine order. Specifically, higher-scale planes are decoded using context obtained from bilinear interpolation of its predecessor, as illustrated in Fig. 2. The distribution parameters for higher-scale features $\mathbf{R}^{l+1}_{c}$ are predicted as:
@@
4.3 Adaptive Quantization
In deep context modeling, each feature is quantized before being processed by the distribution estimation network. However, conventional deep context modeling uses a fixed quantization step size of 1, which is unsuitable for multiscale hexplanes due to their different value range. Additionally, space-only and space-time planes in 4DGS exhibit distinct means and variances owing to differences in spatial and temporal scales as shown in Fig. 3). To address this, we propose an Adaptive Quantization (AQ) method in which different learnable quantization steps $q_{c}$ are assigned to the two groups during 4DGS training. According to [3], quantization can be approximated as adding uniform noise over the quantization range to enable differentiable training. This is illustrated in the total entropy loss for multiscale hexplanes as:
@@ 4.3 Adaptive Quantization
4.4 Entropy-constrained Compression for SH
@@
Figure 4: Qualitative quality comparisons of Standup in the D-NeRF dataset and Flame Steak in the Neu3D dataset. Light4GS achieves up to 65-272× compression with a maximum 1% PSNR degradation. E-D3GS is excluded from the D-NeRF comparison as it is designed only for multi-camera setups.
5 Experiment
@@
5.1 Experimental Configuration
Our model is based on the 4DGS framework [43] using PyTorch [34] and is trained on a single NVIDIA RTX 3090 GPU. We have fine-tuned our initialization and optimization parameters based on the configuration outlined in [43]. More hyperparameters will be shown in the Appendix. The implementation of the checkerboard model is based on [4] with a 4x downscaling on the hyperprior, and the entropy bottleneck is adopted from [3]. During training, we vary $\lambda$ from 5e-5 to 5e-2 and adjust the pruning ratio from 0.19 to 0.64 to achieve different rates.
Table 1: Quantitative results on D-NeRF dataset. The best and the second best results are denoted by red and blue.
@@ 5.1 Experimental Configuration
4DGS [43] 33.03 0.98 0.02 15 mins 78 22
Ours 32.56 0.96 0.03 24 mins 89 2.21
32.77 0.97 0.03 24 mins 87 2.69
32.99 0.98 0.02 25 mins 84 3.66
Dataset: For synthetic scene evaluation, we use D-NeRF dataset [35], which is designed for monocular view synthesis with 50-200 frames from a single camera. For real-world scenes, we use the Neu3D [20] and HyperNeRF [32] datasets: Neu3D is captured using 15-20 cameras, while HyperNeRF employs one or two. Data preprocessing and point cloud initialization are conducted as described in [43].
Baselines: We compare our method with recent NeRF-based and GS-based dynamic NVS approaches. [35, 12, 31, 32, 10, 41, 22, 42, 1] are SOTA NeRF-based approaches. In the GS-based category, time-variant methods include [21, 27, 46, 40], while [43, 49, 2, 25] are selected as representations of time-invariant methods. Since these approaches lack entropy-constrained optimization, we additionally train the SOTA entropy-constrained 3DGS method, CompGS [24], across multiple frames to serve as a baseline. During evaluation, some methods are tested on specific datasets, as they are designed exclusively for either monocular or multi-camera setups.
Table 2: Quantitative results on Neu3D dataset, rendering resolution is set to 1024$\times$768.
@@
Real-Time4DGS [46] 32.01 0.014 0.055 9 hours 114 $>1000$
@@
CompGS [24] 29.61 0.077 0.099 $>1$d $45$ $\sim 825$
4DGS [43] 31.72 0.016 0.049 40 mins 34 38
@@
Note: Coffee martini is excluded in our evaluation. CompGS is trained every 30 frames for estimation.
Metrics: We report the quality of rendered images using peak-signal-to-noise ratio (PSNR), (multiscale) structural similarity index (SSIM), structural desimilarity index (D-SSIM), perceptual quality measure LPIPS, rendering fps, training time and storage. Three rate-distortion points are displayed for each dataset due to space limitation; detailed rate-distortion results can be found in the Appendix.
Bitstream: Light4GS consists of five bitstreams: deformable primitives (excluding SH), a deformation MLP, entropy-coded bitstreams for multiscale hexplanes and SH, and the MHCM. The MHCM, which occupies at most 0.77 MB in our setup, uses separate CNNs for hyperprior and different contexts extraction and distribution prediction. To preserve rendering accuracy, both the MHCM and the primitive’s attributes are stored in float32, except that SH is encoded based on predicted probabilities. AC is applied to encode both multiscale hexplanes and the SH.
5.2 Results
According to the quantitative results in Tables 1-3, we observe consistent findings across all datasets that: 1) All NeRF-based methods render at extremely low speeds ($<2\text{fps}$) due to their cumbersome volumetric rendering process. NeRFs using low-rank grid representations [5, 11] perform well in synthetic scenes but struggle with convergence in real-world scenes and require 100–132× more storage than our approach. 2) GS-based methods generally offer improved rendering quality and speed compared to NeRFs; however, time-variant methods [46, 40, 27] come with significantly larger model sizes and still encounter convergence issues. For example, [27] demands 735× more storage than our approach, and training [21, 27, 46] requires over 8 hours. Deformation-based methods [25, 2, 43] demonstrate better convergence with slightly lower quality, but they still require up to 10× more storage than our approach. 3) Our method outperforms the entropy-constrained GS method, CompGS [24], on Neu3D with a 6.3% PSNR increase and a 218× storage reduction. This underscores the importance of effective temporal modeling, as CompGS’s intra-entropy constraint restricts its performance compared to most baselines. To summarize, Light4GS achieves substantial compression gains not only over NeRF-based and time-variant GS methods but also provides over 10× compression compared to deformable GSs and 200×+ over entropy-constrained GS, without sacrificing other performance metrics.
Table 3: Quantitative results on HyperNeRF’s dataset. Rendering resolution is set to 960×540.
@@ 5.2 Results
Nerfies [31] 22.20 0.803 $\sim$ hours $<$ 1 -
HyperNeRF [32] 22.40 0.814 32 hours $<$ 1 -
@@
Figure 5: Bitrate allocation visualization (space-only). (a): hexplane maps; (b) (c): hexplane bitrate maps of low- and high-scales.
@@
Rate-Distortion and Fidelity Upperbound: To capture the combined rate-distortion of STP, MHCM and entropy-constrained SH compression, we train Light4GS models across various pruning ratios with $\lambda$ setups. The final optimal rate-distortion curve is constructed by taking the convex hull of all resulting points, as illustrated by the solid black line in Fig. 6. Fidelity upperbounds, serving as ground truth for rate-distortion performance, are established by training Light4GS at corresponding pruning ratios with $\lambda$ equal to 0.
Figure 7: Effects of STP, the first (a) and last deformed frame (b) is displayed as toy examples.
5.3 Ablation Study
We conduct detailed introspections of Light4GS’s components and rate-distortion primarily on the D-NeRF dataset for efficiency considerations [^1] If not specified, all ablations are conducted with a 51% STP ratio..
@@ 5.3 Ablation Study
Table 4: Ablation study of MHCM. ”Entropy” denotes entropy-constrained training.
@@
$\checkmark$ $\checkmark$ $\checkmark$ -0.063% / +2.25%
$\checkmark$ $\checkmark$ -0.080% / +8.17%
$\checkmark$ -0.222% / +12.1%
+0.418% / +68.8%
@@
Ablation of MHCM: As shown in Table 4, integrating all deep contexts of MHCM, i.e., checkerboard, inter-scale, and inter-plane, increases the rate-distortion performance. Inter-plane context has the minimal effect, as it applies only to space-only planes at the lowest-scale. The inter-scale context, however, achieves the largest storage savings by eliminating the storage need for high-scale hyperpriors. Although the checkerboard context provides limited storage savings, because the hexplane rate reduction does not offset the increased network size required for context modeling, it has the greatest impact on quality, yielding a 0.14% PSNR increase.
Ablation of Rate-distortion: Table 5 demonstrates the effects of various rate-distortion configurations. Results indicate that applying STP alone achieves a 19.2% reduction in storage with a PSNR drop of 0.43%. Adding the rate-distortion constraint for SH increases storage savings to 35.2%, with PSNR drop of 0.60%. The rate-distortion constraint on multiscale hexplane provides even greater storage reductions than those for SH, as SH coefficients are encoded without applying any deep context. Our approach achieves an 83% storage reduction for primitives and a 98% reduction for multiscale hexplane, while resulting in a loss of PSNR at only 8.47%. These findings emphasize how Light4GS effectively balances significant storage savings with high fidelity through the strategic integration of STP and MHCM.
Impact of Hexplane Features: To explore the impact of scales $l$ and the hexplane feature size $h$, we train Light4GS models with different values of $h$ and $l$. The black anchors in Fig. 8 represent the trained rate-distortion curve based on our experimental setting of $h=32$ and $l=2$. For clearer insights, we conduct ablations for the two parameters at different $\lambda$s. Results suggest that our configuration is near the Pareto-optimal region, where increasing the number of scales yields diminishing PSNR gains since higher-scales struggle to capture unique details beyond those at lower-scales, while reducing scales rapidly degrades the rendering quality. A similar trend holds for feature size ablation.
Table 5: Ablation results showing RD-rates, “GS” denotes Gaussian primitives, “Hex” denotes multiscale hexplanes.
@@
w/o STP &RD $0$%
STP w/o RD -0.43%/-19.2% -43.9% 0%
STP w/ SH-RD, $c_{i}$0.60%/-35.2% -81.2% 0%
STP w/ Hex-RD, $\mathbf{R}_{i}$0.66%/-75.4% 56.1% -98.7%
STP w/ Full-RD, $\mathbf{R}_{i},c_{i}$0.$8$3.1% -98.7%
@@
5.4 Coding Complexity

### v2

@@ IV-C1 Lowest-Scale Hexplane Compression
Planes in one hexplane can be divided into two groups based on their dimensions: space-only and space-time. Throughout the lowest-scale encoding process, both groups employ checkerboard modeling but use different hyperpriors based on the availability of prior information. In this approach, we designate odd positions on each plane as anchors, which are encoded with only the hyperprior. In contrast, non-anchors (even positions) are encoded using both the hyperprior and the checkerboard context that is derived by applying masked convolution on the plane anchors. Without loss of generality, we omit the superscript $l$ (i.e., scaling factor in Eq. (5)) in this section, as it is always equal to 1. To compress the space-time group, since no prior information is available, we directly use a deep network $h_{a}()$ to extract hyperpriors as $\mathbf{z}_{c}=h_{a}(\mathbf{R}_{c})$. The hyperprior is then applied to encode space-time planes as:
@@
IV-C2 High-Scale Hexplane Compression
High-scale hexplanes often exhibit significant redundancy with their lower-scale counterparts. Relying solely on the checkerboard model for compression overlooks these inter-scale redundancies, as it captures only intra-scale correlations. To address this issue, we introduce an inter-scale context that encodes higher-scale planes in a coarse-to-fine order. Specifically, higher-scale planes are decoded using context obtained from bilinear interpolation of its predecessor, as illustrated in Fig. 2. The distribution parameters for higher-scale features $\mathbf{R}^{l+1}_{c}$ are predicted as:
@@ IV-C2 High-Scale Hexplane Compression
Figure 5: Qualitative quality comparisons of Standup in the Neu3D dataset. Light4GS achieves up to 65-272× compression with a maximum 1% PSNR degradation.
IV-D Adaptive Quantization
In deep context modeling, each feature is quantized before being processed by the distribution estimation network. However, conventional deep context modeling uses a fixed quantization step size of 1, which is unsuitable for multiscale hexplanes due to their different value range. Additionally, space-only and space-time planes in 4DGS exhibit distinct means and variances owing to differences in spatial and temporal scales as shown in Fig. 4. To address this, we propose an Adaptive Quantization (AQ) method in which different learnable quantization steps $q_{c}$ are assigned to the two groups during 4DGS training. According to [3], quantization can be approximated as adding uniform noise over the quantization range to enable differentiable training. This is illustrated in the total entropy loss for multiscale hexplanes as:
@@ IV-D Adaptive Quantization
IV-E Entropy-constrained Compression for SH
@@
V Experiment
@@
V-A Experimental Configuration
Our model is based on the 4DGS framework [54] using PyTorch [45] and is trained on a single NVIDIA RTX 3090 GPU. We have fine-tuned our initialization and optimization parameters based on the configuration outlined in [54]. The implementation of the checkerboard model is based on [6] with a 4x downscaling on the hyperprior, and the entropy model is adopted from [3]. During training, we vary $\lambda$ from 5e-5 to 5e-2 and adjust the pruning ratio from 0.19 to 0.64 to achieve different rates. PCA parameters and the hexplane bounds are computed once at the beginning of training from the sparse point cloud.
STP implementation: We implement STP based on [12], where the significance of each primitive is determined by counting the number of intersected pixels. For non-monocular settings, we randomly select a single camera to estimate primitive significance, as evaluating significance across all cameras is computationally expensive. Notably, STP is applied only after the cloning process for primitives has been completed. To avoid abrupt quality degradation, we perform pruning in two passes: one in the middle of training and another near the end, followed by a few fine-tuning iterations.
MHCM Implementation: We build our deep context model on the CompressAI framework [6]. The hyperprior encoder $h_{a}(\cdot)$, decoder $h_{s}(\cdot)$, and entropy model follow the configuration in [3], with a hyperprior that applies $4\times$ downsampling. To extract multiscale deep context, we use a lightweight CNN, where each layer outputs twice as many channels as its input. The checkerboard design with a multiplexer is adopted to accelerate both encoding and decoding [19]. All hexplane dimensions are chosen as multiples of 4 to ensure compatibility with masked convolutions.
TABLE I: Quantitative results on D-NeRF dataset. The best and the second best results are denoted by dark red and light red.
@@ V-A Experimental Configuration
4DGS-HUST [54] 33.03 0.98 0.02 15 mins 78 22
Ours 32.58 0.96 0.03 24 mins 86 2.20
32.79 0.97 0.03 24 mins 85 2.70
33.01 0.98 0.02 25 mins 83 3.67
Dataset: For synthetic scene evaluation, we use D-NeRF dataset [46], which is designed for monocular view synthesis with 50-200 frames from a single camera. For real-world scenes, we use the Neu3D, [29] NeRF-DS [59] and HyperNeRF [43] datasets: Neu3D is captured using 15-20 cameras, while HyperNeRF and NeRF-DS employ one or two. Data preprocessing and point cloud initialization are conducted as described in [54].
Baselines: We compare our method with recent NeRF-based and GS-based dynamic NVS approaches. [46, 16, 42, 43, 14, 53, 31, 52, 1] are SOTA NeRF-based approaches. In the GS-based category, methods such as [30, 37, 57, 51] are time-variant, whereas [54, 64, 2, 35] are time-invariant representations. Since these approaches lack entropy-constrained optimization, we additionally train the latest entropy-constrained 3DGS method, CompGS [34], across multiple frames to serve as a baseline. During evaluation, some methods are tested on specific datasets, as they are designed exclusively for either monocular or multi-camera setups.
[…]
@@
Real-Time4DGS [57] 32.01 0.014 0.055 9 hours 114 >1000>1000
@@
CompGS [34] 29.61 0.077 0.099 >1>1d 45 $\sim 825$
4DGS-HUST [54] 31.72 0.016 0.049 40 mins 34 38
SaroGS [58] 32.03 0.0142 0.044 >>2 hours 40 385
4DGS-FDU [63] 31.57 0.016 0.057 $\sim$9 hours 96 3128
4DGC [21] 31.58 - - - 168 150
MEGA [68] 31.49 0.016 0.056 - 77 25
Compact3DGS [27] 31.73 - 0.053 - 186 22
@@
Note: Test excluding Coffee Martini. CompGS is trained every 30 frames for estimation.
Metrics: We report the quality of rendered images using peak-signal-to-noise ratio (PSNR), (multiscale) structural similarity index (SSIM), structural desimilarity index (D-SSIM), perceptual quality measure LPIPS, rendering frames per second (FPS), training time and storage. Three rate-distortion points are displayed for each dataset due to space limitation.
V-B Training Process
D-NeRF [46]: During the first 3000 iterations, only the standard training process of vanilla 3DGS is applied without any additional techniques. From iterations 3000 to 30000, STP is performed twice based on a predefined ratio, and primitives are densified every 3000 iterations until 14000 iterations. The entropy constraints of the MHCM and SH are applied every 3 iterations, with the SH constraint activated only after densification ends. Adaptive quantization starts at 15000 iterations, with the SH quantization fixed at 0.02 based on pre-evaluation of primitives generated from vanilla 4DGS.
Neu3D, HyperNeRF and NeRF-DS [29, 43, 59]: The training pipeline for the two real-world datasets follows that of D-NeRF, with the only modification being a reduction in training iterations to 20,000.
TABLE III: Quantitative results on NeRF-DS’s dataset. Rendering resolution is set to 960×540.
@@ V-B Training Process
3DGS [24] 20.93 0.7787 - 5 mins 13
D3DGS[64] 23.61 0.8395 35 - 53
MotionGS [17] 24.54 0.8656 35 - 62
4DGS-HUST [16] 23.57 0.8399 46 15 mins 24
Ours 23.59 0.8367 61 20 mins 2.36
[…]
Nerfies [42] 22.20 0.803 $\sim$ hours << 1 -
HyperNeRF [43] 22.40 0.814 32 hours << 1 -
@@
Figure 7: Rate–distortion performance comparison on D-NeRF. Green anchors denote STP only.
Figure 8: Rate–distortion performance comparison on Neu3D.
Figure 9: Rate–distortion performance comparison on HyperNeRF.
Figure 10: Rate–distortion performance comparison on NeRF-DS.
V-C Results
Main Results: According to the quantitative results in Tables I-IV, we observe consistent findings across all datasets that: 1) All NeRF-based methods render at extremely low speeds ($2\text{FPS}$) due to their cumbersome volumetric rendering process. NeRFs using low-rank grid representations [7, 15] perform well in synthetic scenes but struggle with convergence in real-world scenes and require 100–132× more storage than our approach. 2) GS-based methods generally offer improved rendering quality and speed compared to NeRFs; however, time-variant methods [57, 51, 37] come with significantly larger model sizes and still encounter convergence issues. For example, [37] demands 735× more storage than our approach, and training [30, 37, 57] requires over 8 hours. Deformation-based methods [35, 2, 54] demonstrate better convergence with slightly lower quality, but they still require up to 10× more storage than our approach. 3) Our method outperforms the entropy-constrained GS method, CompGS [34], on Neu3D with a 6.3% PSNR increase and a 218× storage reduction. This underscores the importance of effective temporal modeling, as CompGS’s intra-entropy constraint restricts its performance compared to most baselines. (4) On the NeRF-DS dataset, our method achieves a strong compression ratio while maintaining comparable quality. Compared with the 4DGS-HUST anchor, our model reduces storage to 2.36 MB, i.e., about $10\times$ compression, with negligible quality change. Since NeRF-DS scenes are longer and exhibit more complex motion, these results further validate the effectiveness of our approach on both short and long real-world sequences. We also provide a qualitative comparison of Light4GS against existing baselines on the Neu3D dataset, where our method achieves the best rate-distortion performance. Please refer to Fig. 5. To summarize, Light4GS achieves substantial compression gains not only over NeRF-based and time-variant GS methods but also provides over 10× compression compared to deformable GSs and 200×+ over entropy-constrained GS, without sacrificing other performance metrics.
Figure 11: Effects of STP, the first (top) and last deformed frame (bottom) is displayed as toy examples.
Rate Distortion: The total rate-distortion of Light4GS is determined by three key components: SH coefficients, multiscale hexplanes, and STP. The distortion introduced by STP is controlled by predefined pruning ratios, while the distortion from SH coefficients and multiscale hexplanes is influenced by the parameters $\lambda$. In the experiments, the pruning ratios are set to [0.19, 0.36, 0.51, 0.64], with $\lambda$ selected from [5e-2, 1e-2, 5e-3, 1e-3, 5e-4]. The final rate-distortion curve is constructed by taking the convex hull of all trained rate-distortion points, as illustrated by the solid black line in Fig. 12. Fidelity upperbounds, serving as ground truth for rate-distortion performance, are established by training Light4GS at corresponding pruning ratios with $\lambda$ equal to 0.
As shown in Fig. 10-10, on D-NeRF, Light4GS achieves over a 10× reduction in storage compared to 4DGS and HexPlane while preserving competitive visual quality. For Neu3D (Fig. 10), Light4GS demonstrates an 11× reduction in model size relative to 4DGS and achieves an impressive 48× storage reduction over the time-variant 3DGS STG [30]. For HyperNeRF (Fig. 10), Light4GS achieves a 16× storage reduction compared to 4DGS and a 15× reduction relative to per-frame training 3DGS [24]. In complex scenes, Light4GS presents notable benefits as the requirement for additional scales to capture details increases inter-scale similarities, thus improving compression efficiency.
@@ V-C Results
V-D Ablation Study
We conduct detailed introspections of Light4GS’s components and rate-distortion primarily on the HyperNeRF dataset.  [^1] If not specified, all ablations are conducted with a 51% STP ratio..
TABLE V: Ablation study of MHCM. ”Entropy” denotes entropy-constrained training.
@@ V-D Ablation Study
$\checkmark$ $\checkmark$ $\checkmark$ -0.067% / +2.31%
$\checkmark$ $\checkmark$ -0.089% / +8.05%
$\checkmark$ -0.215% / +12.4%
+0.426% / +69.2%
@@
Ablation of MHCM: As shown in Table V, integrating all deep contexts of MHCM, i.e., checkerboard, inter-scale, and inter-plane, increases the rate-distortion performance. Inter-plane context has the minimal effect, as it applies only to space-only planes at the lowest-scale. The inter-scale context, however, achieves the largest storage savings by eliminating the storage need for high-scale hyperpriors. Although the checkerboard context provides limited storage savings, because the hexplane rate reduction does not offset the increased network size required for context modeling, it has the greatest impact on quality, yielding a 0.14% PSNR increase.
Ablation of rate-distortion: Table VI reports an ablation of our MHCM components, where the first row (all enabled) is taken as the reference. Disabling the inter-plane coupling leads to a $2.31\%$ size increase with a $0.067\%$ PSNR drop, and further removing the inter-scale context enlarges the size to $8.05\%$ with a $0.089\%$ PSNR drop. When only entropy-constrained training is kept (i.e., without checkerboard / inter-scale / inter-plane), the size increases to $12.4\%$ and PSNR decreases by $0.215\%$, indicating that the proposed context modules are important for compactness. Finally, removing entropy-constrained training causes a dramatic size increase ($69.2\%$), despite a slight PSNR gain ($0.426\%$), highlighting that entropy-constrained training is the key factor for achieving a favorable rate–distortion trade-off, while checkerboard and cross-scale/plane contexts further improve compression efficiency.
@@
Compressing SH using alternative entropy models: Since a purely fully factorized model may leave bits on the table, we also explore three alternative designs for SH compression: (i) a channel-wise autoregressive model, where the first three SH channels are encoded with a fully factorized entropy model and an MLP predicts the mean and scale of each subsequent channel conditioned on these three; (ii) a coupled context model, where primitive coordinates are used to query hexplane features and the resulting hidden vector is fed into an MLP to predict the conditional Gaussian parameters (mean and scale) shared by all SH channels; and (iii) a mixture-of-logistics model, where each channel is modeled as a mixture of several logistic components rather than a single one, using an 8-dimensional parameter vector per channel while keeping all other settings identical to the fully factorized baseline. From the R-D curves on bouncingballs (Fig. 14), we observe that only the autoregressive model slightly outperforms the fully factorized baseline at higher bitrates, indicating that explicitly modeling correlations between neighboring SH channels can help at the cost of significant coding delay. The coupled context brings no clear gain, as all channels share a single hexplane feature, limiting its ability to capture channel-wise dependencies and coupling GS decoding with hexplane decoding so that the two can no longer be decoded independently or in parallel. The mixture-of-logistics model, while more expressive in principle, likewise shows no noticeable improvement over the single-logistic fully factorized model in our setting. Considering both compression performance and complexity, we therefore use the fully factorized model for SH in our final implementation.
Figure 14: Fully factorized model vs. other entropy models.
TABLE VI: Ablation study on PCA, nonlinear contraction, and rate–distortion on HyperNeRF dataset. “GS” denotes Gaussian primitives; “Hex” denotes multiscale hexplanes.
[…]
Base 0%/0% 0% 0%
w/o PCA -1.31%/- - -
w/o No. Contract 1.23%/- - -
w/ STP -0.35%/-15.9% -48.3% 0%
w/ STP & $c_{i}$-RD -0.63%/-36.5% -86.5% 0%
w/ STP & $\mathbf{R}_{i}$-RD -0.70%/-72.0% 62.4% -95.3%
w/ STP & $\mathbf{R}_{i},c_{i}$-RD -0.90%/-84.5% -88.0% -95.0%
Figure 15: Bitstream composition on Neu3D and DyNeRF (in MB).
@@
V-E Bitstream Composition
Light4GS encodes the scene into five main bitstreams: deformable Gaussian primitives (excluding SH), a deformation MLP, entropy-coded multiscale hexplanes, spherical harmonics (SH), and the MHCM module. The MHCM, occupying no more than 0.77 MB in our setup, employs separate CNNs for hyperprior and context extraction, along with distribution prediction. To maintain rendering fidelity, both the MHCM and Gaussian attributes are stored in Float32, with SH compressed via AC based on predicted probabilities.
Fig. 15 summarizes the bitstream composition across all test scenes, categorized into four components: deformation (MLP), context (MHCM including CNNs and entropy model), hexplane (multiscale features), and GS (Gaussian primitives). Among these, the GS component dominates due to the absence of context modeling in SH compression. The context stream reflects the entire hexplane modeling pipeline, encompassing convolutional, MLP, and entropy coding stages. The relative sizes of deformation and hexplane streams vary with scene complexity, with hexplanes occupying a larger portion in scenes requiring finer spatial detail.
V-F Coding Complexity